OpenStudy (anonymous):

Mallory cut a 6ft long board into 3/4ft pieces. How many 3/4ft pieces did she have ?

OpenStudy (anonymous):

Cutting a board should sound to you like "dividing a board"... that's a hint that this is just a division problem.

OpenStudy (anonymous):

If you divided a 6 foot board into 1 foot pieces, how many pieces would you get? 6 ft divided by 1 ft per piece = 6 pieces

OpenStudy (anonymous):

thanks JakeV8.

OpenStudy (anonymous):

So now you need to do: 6 feet divided by 3/4 foot per piece = 6 / (3/4)

OpenStudy (anonymous):

Since the pieces are shorter, there should be more of them... 6 / (3/4) = 6(4/3) = 24/3 = 8

OpenStudy (anonymous):

Does that make sense? You just have to "translate" the word problem from English to "math language" :)

OpenStudy (anonymous):

thank you so much! it did make sense.

OpenStudy (anonymous):

great, glad to help! :)